OpenStudy (anonymous):

what is the product of 100 g and 18.4cm

Parth (parthkohli):

You can never multiply two different units.

OpenStudy (anonymous):

so how can i multiply them?

Parth (parthkohli):

You can't.

OpenStudy (anonymous):

can't it be the unit is g-cm ?

Parth (parthkohli):

I just realized that. Yes.

Parth (parthkohli):

\[100 g\times 18.4 cm \implies (100 \times 18.4) \times (g \times cm) \implies 1840 \text{gcm} \]

OpenStudy (anonymous):

thanks!